OpenStudy (anonymous):

Can someone PLEASE help me? Will fan/give medal A Labrador leaps over a hurdle. The function f(t) represents the height of the Labrador above the ground, in inches, at t seconds: f(t) = -16t2 + 26t A foxhound jumps over the same hurdle. The table shows the height of the foxhound above the ground g(t), in inches, at t seconds: Time (t) g(t) 0 0 0.4 6.24 0.6 7.44 0.7 7.56 0.8 7.36 1.0 6 1.4 0

OpenStudy (anonymous):

Part A: Compare and interpret the maximum of f(t) and g(t)? (4 points) Part B: Which function has a greater x-intercept? What do the x-intercepts of the graphs of f(t) and g(t) represent? (4 points) Part C: Determine the y-intercepts of both functions, and explain what this means in the context of the problem. (2 points)

OpenStudy (gabebae):

https://answers.yahoo.com/question/index?qid=20140414130742AAFPTrv this might help I took algebra last year early in the year I don't remember how to do most of this but I took it for sure

OpenStudy (kmeis002):

Notice since our quadratic has a \(-a\) term, it opens down and therfore has a maximum halfway between the two zeros. Solving for the zeros we get: \[-16t^2 +26 t = 0\\ 2t(-8t+13)=0 \\2t = 0 \text{ and } -8t+13 = 0\\ t_1= 0 \text{ and } t_2 = \frac{13}{8} \] Finding the halfway point between \( 0 \) and \(\frac{13}{8}\) gives us \(t_{max} = \frac{13}{16} \). So the max output is: \[f(t_{max}) = -16 \left (\frac{13}{16}\right)^2 + 26 \frac{13}{16} = \frac{169}{16} \]

OpenStudy (anonymous):

I thought I had to find the maximum and it would be 13/16 for x?

OpenStudy (anonymous):

Woahhh I think im confusing myself

OpenStudy (anonymous):

By my calculations the maximum is 13/16

OpenStudy (radar):

t is time, and t=13/16 seconds when the function is maximum, have to plug that in to find the maximum height.

OpenStudy (kmeis002):

the maximum occurs at \(t = \frac{13}{16} \), but this is not what the function max

OpenStudy (anonymous):

Ok....So then it would be 169/16?

OpenStudy (kmeis002):

Correct, now you can compare to \(g(t)\)

OpenStudy (anonymous):

So 169/16 which is in decimal form 10.5625 would beat the 0.7 and 7.54?

OpenStudy (anonymous):

0.7 and 7.54 for the foxhound

OpenStudy (kmeis002):

Correct, that max of f is greater than the max of g, what does this mean physically (in terms of the problem)?

OpenStudy (anonymous):

The lab jumps higher?

OpenStudy (kmeis002):

Precisely, so for part b it is askign which one has a greater x-int. Remember x-intercepts occur when \(f(x) = 0\) (or) \(g(x) = 0\). If we can find these, we can compare. I believe g is clear, and the f has already been calculated once in here. What does the x-intercept mean physically?

OpenStudy (kmeis002):

small typo, x =t here.

OpenStudy (anonymous):

The labs x-intercept is (0,0),(138,0)?

OpenStudy (anonymous):

so does that mean either are greater for the intercept?

OpenStudy (anonymous):

neither *

OpenStudy (kmeis002):

excuse me You are correct

OpenStudy (kmeis002):

Yes, so 13/8 was the greatest x int of f(t) which doesn't beat out g(t). What does this tell us physically?

OpenStudy (anonymous):

I'm not sure

OpenStudy (kmeis002):

Ah, so this is measuring the height the dogs jump. That is \(f(t), g(t) \) tells us how high they have reached at some time, t. So if they are both zero, the dogs are where (spatially?)

OpenStudy (anonymous):

they're jumping the same height?

OpenStudy (anonymous):

or one isnt jumping any higher than the other?

OpenStudy (kmeis002):

Well this equation models one jump for each dog. So when you jump you reach a maximum and then you start to fall, eventually you reach the floor. This is when your x-intercept occurs. So which dog hit the ground first?

OpenStudy (anonymous):

The foxhound?

OpenStudy (anonymous):

I'm not sure how much the foxhound jumped because I dont really understand how to map out that kind of table

OpenStudy (kmeis002):

Correct, the x-intercepts for the fox hound are \((0, 0) and (1.4, 0\)). So at t = 1.4, the foxhound hit the floor. The lab hit the floor at t = 1.62. Physically this max sense since the lab jumped higher than the foxhound.

OpenStudy (kmeis002):

Note 1.62 = 13/8

OpenStudy (anonymous):

Oh wow! and thats it for the problem??

OpenStudy (kmeis002):

I have to head out, so finally part c wants to look for the y -ints. This occurs when \(t = 0\). The only y intercepts are \((0,0)\) for both dogs. Physically, this means at time zero they were on the floor since \(f(t), g(t) =0\). This isn't always the case, it is possible for one dog to jump from the ground while the other jumps from a ledge.

OpenStudy (anonymous):

So they would both have the y-intc as (0,0)?

OpenStudy (kmeis002):

Yes, meaning they both started at ground level at the beginning of the jump.

OpenStudy (anonymous):

Ok, thank you so much!

OpenStudy (kmeis002):

You're welcome. I hope it cleared things up.

OpenStudy (anonymous):

I really appreciate it!